OpenStudy (anonymous):

simplify (πd²/4)-(π[d+1]²/4)

OpenStudy (anonymous):

Factor pi/4. Then, it is a difference of squares. Simplify.

OpenStudy (anonymous):

wouldn't it be π/4 (d^2-d^2+1)?

OpenStudy (anonymous):

No. It is pi/4(d^2 - (d+1)^2)

OpenStudy (anonymous):

so how do i solve that? :S

OpenStudy (anonymous):

a^2 - b^2 = (a-b)(a+b)

OpenStudy (anonymous):

In the above, a = d, b = d+1.

OpenStudy (anonymous):

i got π(2d-1)/4 .... is that right? because the back of my book says the answer is π(2d+1)/4

OpenStudy (anonymous):

pi/4 (d^2 - (d+1)^2) = pi/4 (d + d + 1)(d - (d-1)) = pi/4 (2d+1)(d-d+1) = pi/4 (2d+1)

OpenStudy (anonymous):

how is it pi/4 (d + d + 1)(d - (d-1)) ?

OpenStudy (anonymous):

Forget pi/4 since that is already factored out. Focus on (d^2 - (d+1)^2). Like I said, a^2 - b^2 = (a+b)(a-b). Here, you have d^2 - (d+1)^2 = (d+d+1)(d-(d+1))

OpenStudy (anonymous):

OOOOOH ! im subtracting it.. instead of multiplying (x my bad

OpenStudy (anonymous):

wait but i still get 2d-1 :S

OpenStudy (anonymous):

Actually, you get -(2d+1)

OpenStudy (anonymous):

): it says the answer is pi/4 2d+1

OpenStudy (anonymous):

I see the answer as the same, but with a minus sign. But, may be missing something.

OpenStudy (anonymous):

oh well i'll just keep it like this for now and ask my teacher tomorrow. thanks